# SPRINT Enables Interpretable and Ultra-Fast Virtual Screening against Thousands of Proteomes

### v1

@@ Abstract
Virtual screening of small molecules against protein targets can accelerate drug discovery and development by predicting drug-target interactions (DTIs). However, structure-based methods like molecular docking are too slow to allow for broad proteome-scale screens, limiting their application in screening for off-target effects or new molecular mechanisms. Recently, vector-based methods using protein language models (PLMs) have emerged as a complementary approach that bypasses explicit 3D structure modeling. Here, we develop SPRINT, a vector-based approach for screening entire chemical libraries against whole proteomes for DTIs and novel mechanisms of action. SPRINT improves on prior work by using a self-attention based architecture and structure-aware PLMs to learn drug-target co-embeddings for binder prediction, search, and retrieval. SPRINT achieves SOTA enrichment factors in virtual screening on LIT-PCBA and DTI classification benchmarks, while providing interpretability in the form of residue-level attention maps. In addition to being both accurate and interpretable, SPRINT is ultra-fast: querying the whole human proteome against the ENAMINE Real Database (6.7B drugs) for the 100 most likely binders per protein takes 16 minutes. SPRINT promises to enable virtual screening at an unprecedented scale, opening up new opportunities for in silico drug repurposing and development. SPRINT is available on the web as ColabScreen: https://bit.ly/colab-screen
Introduction
@@
Virtual screening has emerged as a powerful tool for predicting drug-target interactions (DTIs) and guiding experimental efforts, but conventional structure-based methods like molecular docking are often too slow for proteome-scale analyses [1]. This limitation hinders their application in crucial parts of the drug discovery process such as off-target prediction [2]. The need for scalable and interpretable virtual screening methods is particularly evident in, for example, antimicrobial drug discovery. The rapid emergence of antimicrobial-resistant pathogens poses a severe threat to public health [3], necessitating the development of new antibiotics with novel mechanisms of action to combat cross-resistances [4]. Effective antimicrobial virtual screening demands methods that are not only fast and scalable but also interpretable, enabling researchers to: 1) identify new drug candidates with on-target effects across thousands of microbial proteomes and minimal off-target effects in humans, and 2) provide interpretations for predicted DTIs and potential mechanisms of action.
@@
Recently, vector-based virtual screening has been proposed as an alternative to structure-based screening to efficiently predict DTIs, leveraging vector featurizations for molecules [5] and sequence models for protein targets [6, 7, 8]. One method, ConPLex [2], proposes co-embedding molecules and proteins into a shared vector space, where the distance between entities is proportional to interaction likelihood. This effectively reduces the task of computing a DTI to a dot product in the co-embedding space, enabling the screening of millions of molecules against the entire human proteome in 24 hours. However, ConPLex does not scale favorably when identifying DTIs across thousands of bacterial and fungal proteomes, and it cannot provide explanations of its DTI predictions. Similarly, DrugCLIP [9, 10] aligns the embeddings of protein pocket structures and ligands with contrastive learning such that similarity encodes the probability of interaction. They demonstrate state-of-the-art (SOTA) results on virtual screening benchmarks with a fraction of the compute time needed for other structure-based virtual screening methods. Their approach is restricted to structures in which binding pockets can be predicted through pocket-detection algorithms or homology-based approaches; however, [11] estimated that almost half of all structured domains may lack obvious pockets in their experimental structures.
@@
In this work, we propose SPRINT (Structure-aware PRotein ligand INTeraction) for fast and accurate vector-based DTI predictions. SPRINT featurizes proteins using SaProt [8], a transformer model trained by augmenting the amino acid vocabulary with discrete structure-tokens [12]. Rather than featurizing proteins by averaging per-residue embeddings from protein sequence models, SPRINT uses a multi-head attention pooling scheme to learn a sequence-dependent aggregation. SPRINT is extremely fast: querying a single protein target against the ENAMINE REAL (6.7B drug) database and predicting its top-100 binders takes 7ms. Our main contributions are summarized as:
  • We achieve a new SOTA on DTI (Table 1) and virtual screening benchmarks (Table 2).
@@
Methods
Results
Table 1: AUPR on test sets for DTI prediction with co-embedding models across benchmarks (mean $\pm$ std). Train, validation, and test splits for BIOSNAP, BindingDB, and DAVIS are taken from [2]. The MERGED dataset is split by homology (see Appendix C for more details). * indicates that we did not do contrastive training on DUD-E with the ConPLex model.
Model Backbone Pooling BIOSNAP DAVIS BindingDB MERGED
ConPLex [2] ProtBert Avg 0.883 $\pm$ 0.004 0.457 $\pm$ 0.037 0.616 $\pm$ 0.009 0.414 $\pm$ 0.004*
ConPLex-attn* ProtBert Attn 0.904 $\pm$ 0.005 0.493 $\pm$ 0.014 0.672 $\pm$ 0.003 0.448 $\pm$ 0.018
SPRINT-xs (10M) SaProt Attn 0.936 $\pm$ 0.001 0.507 $\pm$ 0.005 0.718 $\pm$ 0.0004 0.481 $\pm$ 0.004
SPRINT-sm (16M) SaProt Attn 0.858 $\pm$ 0.001 0.446 $\pm$ 0.003 0.588 $\pm$ 0.0006 0.526 $\pm$ 0.002
Multi-head attention pooling improves DTI prediction.
A limitation of the ConPLex framework is that it averages the per-residue embeddings obtained from PLMs. As much of the relevant signal for DTIs is located in the binding pocket residues, average pooling is prone to noising the contact map information carefully learned by the PLMs through self-attention [15], particularly in the longer sequence length regime. Retraining the ConPLex model with an attention-based, learned aggregation function [16] achieves SOTA predictive scores for DTIs on most benchmarks (Table 1), even when using the same ProtBert model [6].
To see how the learned aggregation scales with the available training data, we trained a SPRINT model on a huge dataset of DTIs, which we refer to as “MERGED”, [17] combining DTI data from PubChem [18], BindingDB [19], and ChEMBL [20]. In total, the MERGED dataset is comprised of 9,067 unique protein targets and 3,529,822 unique ligands, accounting for 854,118 total positive interactions and 80,681,825 total negative interactions (training details can be found in Appendix C). Our largest model, SPRINT-sm, uses 3-layer MLPs to encode molecules and proteins after multi-head attention pooling, in contrast to SPRINT-xs’s single-layer MLPs. SPRINT-sm exhibits overfitting on the BIOSNAP, BindingDB, and DAVIS datasets but significantly improves performance on the much larger MERGED dataset (Table 1), suggesting that there is value in scaling the SPRINT model size as we increase the amount of training data.
@@ Results
SPRINT-ProtBert (13.4M) 73.4 11.9 11.68 10.19 5.27
SPRINT-sm (16M) 73.4 12.3 15.90 10.78 5.29
@@
LIT-PCBA is a challenging, commonly used virtual screening benchmark that addresses biases in the previously used DUD-E dataset [27] to explicitly enable validation of machine learning models. To evaluate the performance of SPRINT models at virtual screening on LIT-PCBA in the zero-shot setting, we pre-trained the deeper SPRINT-sm (16M) model on the MERGED dataset after removing all protein sequences with $\geq$ 90% sequence homology to the LIT-PCBA set using MMSeqs2 [28]. We see that the structure-aware SPRINT models significantly outperform competitor methods in AUROC, BEDROC (alpha = 0.85), and across all enrichment factor thresholds (Table 2). The SPRINT models outperform similarly sized models trained using ProtBert featurizations and multi-head attention pooling (SPRINT-ProtBert), and models trained using SaProt featurizations and average pooling (SPRINT-Average) demonstrating the importance of structure and self-attention.
Structure-aware protein embeddings improve attention maps.
@@
Following training on the MERGED dataset with either ProtBert or SaProt as the PLM backbone for SPRINT, we analyze the attention patterns learned on a set of single-chain protein-ligand binding structures. We find that the models with the greatest enrichment factors on LIT-PCBA, trained with increased negative sampling, have sparse attention which focuses on residues very distant from ligand interactions (Figures S2, S3, and S5). Therefore, we focus our attention analysis on SPRINT models trained with equal positive and negative sampling (LIT-PCBA results are provided in Table S2). All but one of the ProtBert attention heads attend less to the binding residues than the non-binding residues (Figure 2(a)). By introducing explicit knowledge of the protein’s structure with SaProt, we increase the number of heads attending to the binding residues more than the non-binding residues on average (Figure 2(b)). We visualize the attention on the bound structure of a serine/threonine kinase (PDB ID: 2X4Z) in Figure 3 (additional visualizations provided in Appendix E). Both models have sparse attention maps, with only a handful of residues with non-trivial attention values per head. Attention head 2 of the SaProt model pulls out several residues near the binding site of the kinase while none of the Protbert heads have much, if any, attention on the residues near the binding site. Most of the residues selected by the ProtBert models are on the edges of the protein. We compare the attention patterns to a multiple-sequence alignment (MSA) of 497 human kinase domains from [29] and find that both models attend to non-conserved residues of the kinase which could identify the exact protein with a small residue fingerprint. While the learned aggregation layer allows for model interpretation, we find there is little biological relevance for the attention patterns of the model at its current scale.
(a) ProtBert
(b) SaProt
@@
Figure 2: Comparing the average attention weight of binding and non-binding residues on our set of 109 single-chain protein-ligand binding structures after training on the MERGED Dataset (Methodology detailed in Appendix D). We visualize the Protbert and SaProt models trained with equal positive and negative sampling. The horizontal line indicates the average across the proteins. Visualizations of the ProtBert and SaProt models trained with increased negative sampling are in Figure S2).
ProtBert Refer to caption SaProt Refer to caption
Head 1
Head 2
Head 3
Head 4
@@
Figure 3: Analyzing the attention on PDB ID 2X4Z using ProtBert and SaProt models trained with equal ratio of positive and negative examples (identical models trained with different initial random seeds visualized in Figures S4 and S6; models trained with increased negative sampling visualized in Figures S3 and S5). Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
SPRINT enables querying binding partners from 5132 proteomes.
@@
To demonstrate the utility of SPRINT at the pan-species proteome scale, we constructed a dataset of 5,043 bacterial proteomes, 88 fungal proteomes, and the human proteome, containing 4,291,525 total protein sequences. To store and query the co-embeddings, we use the Chroma vector store[30], a tool developed for semantic search and retrieval-augmented generation in natural language processing. The scaling properties of this framework are highly favorable (Fig. 4): querying a ligand for the 100 most likely binders against the entirety of UniProt (60M sequences) takes 0.0001s, and querying all $62\mathrm{e}$ 6 molecules in CHEMBL for each of their 10 most likely binders against the 4.3M proteins in our multi-species dataset takes less than 4 hours. As a proof of concept, we co-visualized several antimicrobials and drugs with their known protein targets across microbial proteomes (Fig. S1).
Pre-training to predict DTIs improves property prediction.

### v2

@@ Abstract
Virtual screening of small molecules against protein targets can accelerate drug discovery and development by predicting drug-target interactions (DTIs). However, structure-based methods like molecular docking are too slow to allow for broad proteome-scale screens, limiting their application in screening for off-target effects or new molecular mechanisms. Recently, vector-based methods using protein language models (PLMs) have emerged as a complementary approach that bypasses explicit 3D structure modeling. Here, we develop SPRINT, a vector-based approach for screening entire chemical libraries against whole proteomes for DTIs and novel mechanisms of action. SPRINT improves on prior work by using a self-attention based architecture and structure-aware PLMs to learn a co-embedding space for drugs and targets, enabling efficient binder prediction, search, and retrieval. SPRINT achieves SOTA enrichment factors in virtual screening on LIT-PCBA, DTI classification benchmarks, and binding affinity prediction benchmarks, while providing interpretability in the form of residue-level attention maps. In addition to being both accurate and interpretable, SPRINT is ultra-fast: querying the whole human proteome against the ENAMINE Real Database (6.7B drugs) for the 100 most likely binders per protein takes 16 minutes. SPRINT promises to enable virtual screening at an unprecedented scale, opening up new opportunities for in silico drug repurposing and development. SPRINT is available on the web as ColabScreen: https://bit.ly/colab-screen.
keywords:
virtual screening, structure-based drug discovery, protein language models
1 Introduction
@@
Virtual screening has emerged as a powerful tool for predicting drug-target interactions (DTIs) and guiding experimental efforts, but conventional structure-based methods like molecular docking are often too slow for proteome-scale analyses [1]. This limitation hinders their application in crucial parts of the drug discovery process such as off-target prediction [2]. The need for scalable and interpretable virtual screening methods is particularly evident in, for example, antimicrobial drug discovery. The rapid emergence of antimicrobial-resistant pathogens poses a severe threat to public health [3], necessitating the development of new antibiotics with novel mechanisms of action to combat cross-resistances [4]. Effective antimicrobial virtual screening requires methods that are not only fast and scalable but also interpretable, enabling researchers to: 1) identify new drug candidates with on-target effects across thousands of microbial proteomes and minimal off-target effects in humans and 2) provide interpretations for predicted DTIs and potential mechanisms of action.
@@
Recently, vector-based virtual screening has been proposed as an alternative to structure-based screening to efficiently predict DTIs, leveraging vector featurizations for molecules [5] and sequence models for protein targets [6, 7, 8]. One method, ConPLex [2], proposes co-embedding molecules and proteins into a shared vector space, where the distance between entities is proportional to interaction likelihood. This effectively reduces the task of computing a DTI to a dot product in the co-embedding space, enabling the screening of millions of molecules against the entire human proteome in 24 hours. However, ConPLex does not scale favorably when identifying DTIs across thousands of bacterial and fungal proteomes, and it cannot provide explanations of its DTI predictions. Similarly, DrugCLIP [9, 10] aligns the embeddings of protein pocket structures and ligands with contrastive learning such that similarity encodes the probability of interaction. They demonstrate state-of-the-art (SOTA) results on virtual screening benchmarks with a fraction of the compute time needed for other structure-based virtual screening methods. Their approach is restricted to structures for which binding pockets can be predicted using pocket detection algorithms or homology-based approaches; however, [11] estimated that almost half of all structured domains may lack obvious pockets in their experimental structures.
@@
In this work, we propose SPRINT (Structure-aware PRotein ligand INTeraction) for fast and accurate vector-based DTI predictions. SPRINT implicitly uses structure information by featurizing proteins with SaProt [8], a transformer model trained by augmenting the standard amino acid vocabulary with discrete structure-tokens [12]. Instead of averaging the per-residue embeddings from SaProt, SPRINT uses a multi-head attention pooling scheme to learn a sequence-dependent aggregation for protein representation.
SPRINT is extremely fast: querying a single protein target against the ENAMINE REAL (6.7B drugs) database and predicting its top-100 binders takes 7ms when utilizing ChromaDB[13]. Our main contributions are summarized as:
  • Achieving excellent performance on DTI classification (Table 1), virtual screening (Table 2), and binding affinity prediction (Table 3).
  • Enhance the second place method in the second critical assessment of computational hit-finding experiments (CACHE) challenge.
@@
2 Results
Table 1: AUPR on test sets for DTI prediction with co-embedding models across benchmarks (mean $\pm$ std). Train, validation, and test splits for BIOSNAP, Unseen Drugs, Unseen Targets, BindingDB, and DAVIS are taken from [2]. The MERGED dataset is split by homology (see Appendix C for more details). * indicates that we did not do contrastive training on DUD-E with the ConPLex model.
Model ConPLex ConPLex-attn* SPRINT-sm (10M) SPRINT (16M)
Backbone ProtBert ProtBert SaProt SaProt
Pooling Avg Attn Attn Attn
BIOSNAP 0.883$\pm$0.004 0.904$\pm$0.005 0.936$\pm$0.001 0.858$\pm$0.001
Unseen Drugs 0.874$\pm$0.002 0.905$\pm$0.002 0.906$\pm$0.001 0.851$\pm$0.002
Unseen Targets 0.842$\pm$0.006 0.844$\pm$0.011 0.849$\pm$0.006 0.793$\pm$0.007
DAVIS 0.457$\pm$0.037 0.493$\pm$0.014 0.507$\pm$0.005 0.446$\pm$0.003
BindingDB 0.616$\pm$0.009 0.672$\pm$0.003 0.718$\pm$0.0004 0.588$\pm$0.0006
MERGED 0.414$\pm$0.004* 0.448$\pm$0.018 0.481$\pm$0.004 0.526$\pm$0.002
2.1 Multi-head attention pooling improves DTI prediction
A limitation of the ConPLex framework is that it averages the per-residue embeddings obtained from PLMs [2]. As much of the relevant signal for DTIs is located in the binding pocket residues, average pooling can noise the contact map information carefully learned by PLMs through self-attention [14], particularly in the longer sequence-length regime. Retraining the ConPLex model with an attention-based, learned aggregation function [15] achieves SOTA predictive scores for DTIs on most benchmarks (Table 1), even when using the same ProtBert model [6].
To see how the learned aggregation scales with the available training data, we trained a SPRINT model on a huge dataset of DTIs, which we refer to as “MERGED”, [16] combining DTI data from PubChem [17], BindingDB [18], and ChEMBL [19] (further details can be found in Appendix C). Our largest model, SPRINT, uses 3-layer MLPs to encode molecules and proteins after multi-head attention pooling, in contrast to SPRINT-sm’s single-layer MLPs. SPRINT exhibits overfitting on the BIOSNAP, BindingDB, and DAVIS datasets but significantly improves performance on the much larger MERGED dataset (Table 1), confirming that there is value in scaling the SPRINT model size as we increase the amount of training data.
@@ 2.1 Multi-head attention pooling improves DTI prediction
SPRINT-ProtBert (15.9M) 73.4 11.9 11.68 10.19 5.27
SPRINT (16M) 73.4 12.3 15.90 10.78 5.29
@@
LIT-PCBA[26] is a challenging, commonly used virtual screening benchmark that addresses biases in the previously used DUD-E dataset [27] to explicitly enable validation of machine learning models. However, the activity labels in LIT-PCBA are derived from dose-response bioassays reported in PubChem [17], which may introduce noise and variability that could potentially impact the reliability of model evaluations. To evaluate the performance of SPRINT models at virtual screening on LIT-PCBA in the zero-shot setting, we pre-trained the deeper SPRINT (16M) model on the MERGED dataset after removing all protein sequences with $\geq$ 90% sequence homology to the LIT-PCBA set using MMSeqs2 [28]. We see that the structure-aware SPRINT models significantly outperform competitor methods in AUROC, BEDROC (alpha = 0.85), and across all enrichment factor thresholds (Table 2). The SPRINT models outperform similarly sized models trained using ProtBert featurizations and multi-head attention pooling (SPRINT-ProtBert), and models trained using SaProt featurizations and average pooling (SPRINT-Average) demonstrating the importance of structure and self-attention.
We further evaluate SPRINT for binding affinity prediction on the Therapeutic Data Commons (TDC) ‘BindingDB_Patent’ Leaderboard[29]. We see in Table 3 that SPRINT using both ProtBert, ESM2, and SaProt PLMs are competitive with the top-ranking model on the leaderboard, Otter-Knowledge-Ensemble[30]. This is notable, as the top ranking model is an ensemble of four knowledge graph-refined protein and ligand representations with each set of representations fine-tuned on a separate dataset of protein-ligand interactions. All of the models start with ESM2 and Morgan fingerprint embeddings for the protein and ligand, respectively. We see that the SPRINT-ESM2 model performs similar to the Otter-Knowledge-Ensemble, an ensemble of refined ESM2 and Morgan representations. The SPRINT models match the performance while only observing the BindingDB_Patent training set. Interestingly, both ProtBert and ESM2 outperform SaProt in this task, despite SaProt performing better in previous tasks (Tables 1 and 2).
Table 3: Pearson’s Correlation Coefficient on the TDC BindingDB_Patent Leaderboard (mean $\pm$ std).
Model Pearson’s Correlation Coefficient
Otter-Knowledge-Ensemble[30] $0.588\pm 0.002$
SPRINT-ProtBert $0.593\pm 0.005$
SPRINT-ESM2 $0.582\pm 0.012$
SPRINT $0.588\pm 0.011$
2.2 Enhancing Real-World Virtual Screening in CACHE2
The Critical Assessment of Computational Hit-Finding Experiments (CACHE) [31] serves as a benchmark for prospective evaluation of virtual screening methods through experimental validation of predicted hits. CACHE challenge 2 focused on the RNA binding site of SARS-CoV-2 NSP13, a helicase representing the most conserved site across coronaviruses [32] with no known inhibitors. . Crystal structures with bound fragments in the RNA-binding site exist, with PDB ID 5RLZ used for virtual screening. The second-place team [33] implemented a DeepDocking [34] approach using Gnina [1, 23]. They first filtered the Enamine REAL database for drug-like properties, reducing it to 4.4 billion molecules. Then, a random sample of 100,000 molecules (Batch 0) was docked against the 5RLZ structure using Gnina, and a surrogate model was trained to predict the docking score directly from a fingerprint of the molecule [5]. The entire database was scored with the surrogate model, and the top-scoring 100,000 molecules (Batch 1) were docked again with Gnina to produce training data for a new surrogate model. This process was repeated until the docking scores converged (Batch 5), as shown in Figure 1.
To evaluate the utility of SPRINT in filtering a compound database for high-likelihood binders and speeding up the DeepDocking process, we used SPRINT, pretrained on the MERGED dataset, to predict DTI scores for the filtered Enamine REAL database against the same virtual screening structure (5RLZ). We selected the highest-scoring 100,000 molecules from the filtered database and docked them using Gnina, following the same docking protocol as [33]. The virtual screening scores (CNN VS [1]) of our docked compounds are compared to the batches of Gnina DeepDocking in Figure 1. Molecules selected by SPRINT had a higher average CNN VS score (3.10) than Batch 0 of DeepDocking (2.59), and are more distributionally similar to Batches 1, 2, and 3. Notably, SPRINT finds 16 diverse, high-scoring molecules (CNN VS $>$ 6) shown in Figure 7, while only 6 high-scoring molecules are found through DeepDocking (with one high-scoring molecule found during the initial random selection, Batch 0).
Figure 1: Gnina CNN VS scores of the molecules found during DeepDocking and the molecules picked with SPRINT.
2.3 Structure-aware protein embeddings improve attention maps
@@
Following training on the MERGED dataset with either ProtBert or SaProt as the PLM backbone for SPRINT, we analyze the attention patterns learned on a set of single-chain protein-ligand binding structures. We find that models with the greatest enrichment factors on LIT-PCBA, which are trained by sampling many negative DTIs for each positive DTI, have sparse attention that focuses on residues very distant from ligand interactions (Figures 8, 9, and 11). Therefore, we focus our attention analysis on SPRINT models trained with equal positive and negative sampling (LIT-PCBA results are provided in Table 6). All but one of the ProtBert attention heads attend less to the binding residues than they do to the non-binding residues (Figure 2). By introducing explicit knowledge of the protein’s structure with SaProt, we find more heads attending to the binding residues than the non-binding residues on average. We visualize the attention on the bound structure of a serine/threonine kinase (PDB ID: 2X4Z) in Figure 3 (additional visualizations provided in Appendix F). Both models have sparse attention maps, with only a handful of residues with non-trivial attention values per head. Attention head 2 of the SaProt model pulls out several residues near the binding site of the kinase, while none of the Protbert heads have much, if any, attention on the residues near the binding site, instead focusing on the edges of the protein.
Although the learned aggregation layer allows for model interpretation, we find there is little biological relevance for the attention patterns of the model at its current scale. For example, we compare the attention patterns to a multiple sequence alignment (MSA) of 497 human kinase domains from [35] and find that both models attend to non-conserved residues of the kinase that could identify the exact protein with a small residue fingerprint.
@@
Figure 2: Comparing the average attention weight of binding and non-binding residues on our set of 109 single-chain protein-ligand binding structures after training on the MERGED Dataset (Methodology detailed in Appendix E). We visualize the Protbert and SaProt models trained with equal positive and negative sampling. The horizontal line indicates the average across the proteins. Visualizations of the ProtBert and SaProt models trained with increased negative sampling are in Figure 8).
@@
Figure 3: Analyzing the attention on PDB ID 2X4Z using ProtBert and SaProt models trained with equal ratio of positive and negative examples (identical models trained with different initial random seeds visualized in Figures 10 and 12; models trained with increased negative sampling visualized in Figures 9 and 11). Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
2.4 SPRINT enables querying binding partners from 5132 proteomes
@@
To demonstrate the utility of SPRINT at the pan-species-proteome scale, we constructed a dataset of 5,043 bacterial proteomes, 88 fungal proteomes, and the human proteome, containing 4,291,525 total protein sequences. To store and query the co-embeddings, we use the Chroma vector store[13], a tool developed for semantic search and retrieval-augmented generation in natural language processing. The scaling properties of this framework are highly favorable (Fig. 4): querying a ligand for the 100 most likely binders against the entirety of UniProt (60M sequences) takes 0.0001s, and querying all $62\mathrm{e}$ 6 molecules in CHEMBL for each of their 10 most likely binders against the 4.3M proteins in our multi-species dataset takes less than 4 hours. As a proof of concept, we co-visualized several antimicrobials and drugs with their known protein targets across microbial proteomes (Fig. 6).
2.5 Pre-training to predict DTIs improves property prediction